OpenStudy (haleyelizabeth2017):

Describe the end behavior of the graph of the function f(x)=3x^3.

Nnesha (nnesha):

when the highest degree is odD: if the leading coefficient is negative then left end will falls and right end will rise if the leading coefficient is positive then graph rises to the left and right end will falls

OpenStudy (haleyelizabeth2017):

Is it \[x \rightarrow \infty, y \rightarrow \infty ~~and~~ x \rightarrow -\infty,y \rightarrow -\infty \]

Nnesha (nnesha):

looks good

OpenStudy (haleyelizabeth2017):

Woohoo! Thank you!

Nnesha (nnesha):

np